# Decoding Speculative Decoding

### v2

@@ Abstract
Speculative Decoding is a widely used technique to speed up inference for Large Language Models (LLMs) without sacrificing quality. When performing inference, speculative decoding uses a smaller draft model to generate speculative tokens and then uses the target LLM to verify those draft tokens. The speedup provided by speculative decoding heavily depends on the choice of the draft model. In this work, we perform a detailed study comprising over 350 experiments with LLaMa-65B and OPT-66B using speculative decoding and delineate the factors that affect the performance gain provided by speculative decoding. Our experiments indicate that the performance of speculative decoding depends heavily on the latency of the draft model, and the draft model’s capability in language modeling does not correlate strongly with its performance in speculative decoding. Based on these insights we explore a new design space for draft models and design hardware-efficient draft models for speculative decoding. Our newly designed draft model for LLaMa-65B can provide 60% higher throughput than existing draft models and can generalize further to the LLaMa-2 model family and supervised fine-tuned models.
@@ 1 Introduction
  • We show the need for a systematic redesign of draft models used for speculative decoding. We demonstrate that using accuracy on language modeling tasks to choose the draft model for speculative decoding can lead to suboptimal choices, and our experiments highlight that redesigning draft models can improve the throughput of speculative decoding by up to 60%. Based on these insights, our pruned LLaMa-796M provides up to $60\%$ higher throughput than Sheared-LLaMa-1.3B while using only $0.8\%$ of tokens (0.42B vs 50.42B) used to train Sheared-LLaMa-1.3B. We also show that LLaMa-796M works well for other LLMs, such as LLaMa-2 families of models and supervised fine-tuned models (Vicuna-33B).
@@ 3 Understanding Speculative Decoding
We implement speculative decoding in the Microsoft Deepspeed library Microsoft (2023). We use the same setup as SpecInfer Miao et al. (2023), first using the draft model to generate draft tokens and then using the target model to verify the output of the draft model. We set the batchsize to 1 and use greedy decoding. For all our experiments, we use 4 Nvidia 80GB A100 GPUs. We perform our experiment on the OPT and LLaMa base models Zhang et al. (2022); Touvron et al. (2023a) on MMLU Hendrycks et al. (2020), Hellaswag Zellers et al. (2019), and Chatbot Arena datasets Zheng et al. (2023). For MMLU, we use the standard 5-shot setup. The remaining datasets were evaluated in a zero-shot setting. We use OPT-66B and LLaMa-65B as the target LLM for OPT and LLaMa series and use OPT-125M, OPT-350M, OPT-1.3B, OPT-2.7B, and OPT-6.7B as draft models for OPT series, and LLaMa-7B, and LLaMa-13B as draft models for LLaMa series.
@@ 3.3 Understanding Draft Model TAR
In this section, we perform experiments to understand the correlation between the accuracy of a model on popular NLP tasks and its TAR. We plot the accuracy of a model against the TAR it achieves in Figure 5. Surprisingly, we find that TAR has little correlation to the model’s accuracy on a task. We believe this lack of correlation is due to the majority of tokens in a sentence not being content words, which do not affect the accuracy of the model on a specific task.
@@ 4.2 Ablation Studies
In this section, we study if using a different or supervised fine-tuned target model would affect our draft model’s performance.

### v3

@@ Abstract
Speculative Decoding is a widely used technique to speed up inference for Large Language Models (LLMs) without sacrificing quality. When performing inference, speculative decoding uses a smaller draft model to generate speculative tokens and then uses the target LLM to verify those draft tokens. The speedup provided by speculative decoding heavily depends on the choice of the draft model. In this work, we perform a detailed study comprising over 350 experiments with LLaMa-65B and OPT-66B using speculative decoding and delineate the factors that affect the performance gain provided by speculative decoding. Our experiments indicate that the performance of speculative decoding depends heavily on the latency of the draft model, and the draft model’s capability in language modeling does not correlate strongly with its performance in speculative decoding. Based on these insights we explore a new design space for draft models and design hardware-efficient draft models for speculative decoding. Our newly designed draft model for LLaMa-65B can provide 111% higher throughput than existing draft models and can generalize further to the LLaMa-2 model family and supervised fine-tuned models.
@@ 1 Introduction
  • We show the need for a systematic redesign of draft models used for speculative decoding. We demonstrate that using accuracy on language modeling tasks to choose the draft model for speculative decoding can lead to suboptimal choices, and our experiments highlight that redesigning draft models can improve the throughput of speculative decoding by up to 111% in sampling-based decoding and 60% in greedy decoding. Based on these insights, our pruned LLaMa-796M provides up to $111\%$ higher throughput than Sheared-LLaMa-1.3B while using only $0.8\%$ of tokens (0.42B vs 50.42B) used to train Sheared-LLaMa-1.3B. We also show that LLaMa-796M works well for other LLMs, such as LLaMa-2 families of models and supervised fine-tuned models (Vicuna-33B).
@@ 3 Understanding Speculative Decoding
We implement speculative decoding in the Microsoft Deepspeed library Microsoft (2023). We use the same setup as SpecInfer Miao et al. (2023), first using the draft model to generate draft tokens and then using the target model to verify the output of the draft model. We set the batchsize to 1 and use greedy decoding. For all our experiments, we use 4 Nvidia 80GB A100 GPUs. We perform our experiment on the OPT and LLaMa base models Zhang et al. (2022); Touvron et al. (2023a) on MMLU Hendrycks et al. (2020), Hellaswag Zellers et al. (2019), and Chatbot Arena datasets Zheng et al. (2023). For MMLU, we use the standard 5-shot setup. The remaining datasets were evaluated in a zero-shot setting. Note that since our goal is to test our draft model’s ability to study the target model’s behavior, we do not instruct the model to emit a single-letter answer to MMLU and Hellaswag questions, but instead opt for an open-ended generation approach where the model can provide as much as explanation as it sees fit, which aligns better with a real-world chatbot setting. For MMLU, we feed the model with the question and choices; for Hellaswag, we feed the model with incomplete sentences without the choices. We use OPT-66B and LLaMa-65B as the target LLM for OPT and LLaMa series and use OPT-125M, OPT-350M, OPT-1.3B, OPT-2.7B, and OPT-6.7B as draft models for OPT series, and LLaMa-7B, and LLaMa-13B as draft models for LLaMa series.
@@ 3.3 Understanding Draft Model TAR
In this section, we perform experiments to understand the correlation between the accuracy of a model on popular NLP tasks and its TAR. We plot the accuracy of a model against the TAR it achieves in Figure 5. Surprisingly, we find that TAR correlates little to the model’s accuracy on a task. We believe this lack of correlation is due to the majority of tokens in a sentence not being content words Chen et al. (2023), which do not affect the model’s accuracy on a specific task.
@@ 4.1 Draft Model Design
Furthuer, our model also reduces intermediate memory usage by $37\%$ compared to the original model. Our memory saving mainly comes from wider MLP layers, which does not require storing KV-caches.
@@ 4.2 Ablation Studies
In this section, we study if varying the decoding sampling methods or using a different or supervised fine-tuned target model would affect our draft model’s performance.
Temperature sampling: Next, we perform ablation by varying the sampling method. In our prior experiments, we used the greedy decoding method. However, sampling-based decoding methods often improve LLM decoding quality Holtzman et al. (2019); Ficler and Goldberg (2017); Fan et al. (2018). To understand the effect of sampling on our findings, we performed experiments with temperature sampling by choosing temperatures of 0.5 and 1. In Table 3, we show that our model achieves even higher speedup than existing models when stochasticity is introduced to the decoding process. On three datasets we achieved from $97\%$ to $111\%$ higher throughput using our NoFT-Wide-796M model compared to using an existing fine-tuned model. The speedups in different temperature settings show that our speedups generalize across different sampling methods.
Table 3: This table shows the throughput of speculative decoding (tokens/s) with temperature sampling on various datasets and models. Our NoFT-Wide-796M model achieved $97.7\%$ to $111.8\%$ higher throughput comparing to the existing Sheared-LLaMA-1.3B model while only using $0.8\%$ of its training tokens.
Temperature = 1.0 Temperature = 0.5
Draft Model MMLU Hellaswag Chatbot Arena MMLU Hellaswag Chatbot Arena
NoFT-Wide-543M 20.54 24.08 23.77 26.08 25.77 25.31
NoFT-Wide-796M 24.32 24.32 24.18 29.64 26.91 25.76
NoFT-Wide-1.3B 22.99 24.08 24.22 28.43 26.39 25.85
Sheared-LLaMA-1.3B 12.06 12.30 11.45 14.60 13.45 12.66
Speedup 101.6% 97.7% 111.8% 103.0% 100.0% 103.5%